OpenStudy (anonymous):

Give exact and approximate solutions to three decimal places y^2 -4+4=9 y=

OpenStudy (anonymous):

\[y^2 -4+4=9\]

OpenStudy (anonymous):

-+3

OpenStudy (anonymous):

I'm guessing this is supposed to be y^2 -4x+4=9 -< (y-2)^2 = 9 take square root

OpenStudy (sasogeek):

y^2-0y-9=0

OpenStudy (sasogeek):

solve the quadratic, get ur answer

OpenStudy (anonymous):

That's how to find the solution?

OpenStudy (anonymous):

To repeat I'm guessing this is supposed to be y^2 -4x+4=9 (y-2)^2 = 9 take square root

OpenStudy (sasogeek):

y^2-4+4=9 -4+4=0 => y^2=9 y=root of 9 y=3

OpenStudy (anonymous):

so it's 5, -1

OpenStudy (anonymous):

Yes.

OpenStudy (sasogeek):

to 3 decimal places....

OpenStudy (anonymous):

I dont understand that part

OpenStudy (sasogeek):

me neither

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

maybe I square each solution... idk